OpenStudy (anonymous):

I REALLY NEED HELP a store sells cashews for $4.00 per pound and peanuts for $1.50 per pound. the manager decides to mix 20 pounds of peanuts with some cashews and sell the mixture for $2.00 per pound. how many pounds of cashews should be mixed with the peanuts so that the mixture will produce the same revenue as would selling the nuts separately?

OpenStudy (anonymous):

if u have any ideas on how to do this it would be greatly appreciated

OpenStudy (anonymous):

please

OpenStudy (anonymous):

(pounds of peanuts)(price of peanuts per pound) + (pounds of cashews)(price of cashews)=(totalpounds)(price of pounds) so then 20(1.5) + x(4) = (20+x)2 so x=5 it should be 5 pounds of cashews